OpenStudy (anonymous):

How do you put it in polynomial from f(x)=-1/2(x+4)^2-5

OpenStudy (anonymous):

The expression is not poly

OpenStudy (anonymous):

i know but how do you put it in polynomial form

OpenStudy (athe):

Is it right? Is it you equation? \[f(x)=\frac{1}{2}(x+4)^2-5\]

OpenStudy (anonymous):

hello you can write -1/2 (x^2+16+8x)-5 =- 1/2x^2-8-4x-5 polynomial form is - 1/2x^2-4x-13